OpenStudy (anonymous):

y=4x-7 , y=2x+9 how can i solve this by using substitution

OpenStudy (anonymous):

When you see an equals sign, it means that the thing on the left is the EXACT SAME THING as the thing as the right. Whenever you see the thing on the left, you can replace it with the thing on the right... because they're the same thing. Substitution uses this to make an easier problem. In this example, you must solve for two variables, x and y. Using substitution, you can solve for just one, and then substituting again, find the other. y = 4x - 7 y = 2x + 9 Replacing first y with 2x + 9 2x + 9 = 4x - 7 Subtract 9. 2x = 4x - 16 Subtract 4x. -2x = -16 Divide by -2/ x = 8 Substituting the new value of x in the first equation, y = 4x - 7 y = 4*8 - 7 y = 32 - 7 = 25 x = 8, y = 25 (8, 25) Plug the answer back in to make sure everything works.

OpenStudy (anonymous):